OpenStudy (anonymous):

Find the rate of change : (0,9) and (-3, -4

jimthompson5910 (jim_thompson5910):

This is the same as finding the slope of the line through these two points. So we use the slope formula. m = (y2-y1)/(x2-x1) m = (-4-9)/(-3-0) m = (-13)/(-3) m = 13/3 So the slope of the line that passes through the two points is 13/3 which means that the rate of change is 13/3